OpenStudy (anonymous):

How do I write a recursive rule for the geometric series -5, -15, -45, -135?

OpenStudy (anonymous):

\[-5,-5\times 3,-5\times 3^2,...\]

OpenStudy (anonymous):

i guess as a recursion you could say \(a_1=-5,a_n=a_{n-1}\times 3\)

OpenStudy (marigirl):

we can use the formula xn = ar^(n-1) where a is the first term r is the 'common ratio'

OpenStudy (marigirl):

as satellite73 pointed out -5 -15 -45 -135 −5×3 −5×3^2 -5×3^3 (difference) from that we can conclude the common ration r=3 .... hope that makes sense. happy to clarify further

OpenStudy (marigirl):

what do u conclude as your geometrical seq rule for this?

OpenStudy (anonymous):

a(1) = -5 a(n+1) = 5 * a(n)

OpenStudy (anonymous):

i mean a(n+1) = 3 * a(n)

OpenStudy (anonymous):

I understand the common ratio and the first term but the n-1 is throwing me off, that's why I'm not sure how to write it, I already figured out the common ratio and the iterative rule for the sequence but I just don't get the recursive rule, none of this makes any sense to me at all.

OpenStudy (anonymous):

follow on IG to get in conact eith me MARTIZaboo

OpenStudy (anonymous):

I don't go on IG.

OpenStudy (anonymous):

okay cool